### v2

# Algorithmic Information Theory for Graph Edge Grouping and Substructure Analysis

Abstract
Understanding natural phenomenon through the interactions of different complex systems has become an increasing focus in scientific inquiry. Defining complexity and actually measuring it is an ongoing debate and no standard framework has been established that is both theoretically sound and computationally practical to use. Currently, one of the fields which attempts to formally define complexity is in the realm of Algorithmic Information Theory. The field has shown advances by studying the complexity values of binary strings and 2-dimensional binary matrices using 1-dimensional and 2-dimensional Turing machines, respectively. Using these complexity values, an algorithm called the Block Decomposition Method developed by Zenil, et al. in 2018, has been created to approximate the complexity of adjacency matrices of graphs which have found relative success in grouping graphs based on their complexity values. We use this method along with another method called edge perturbation to exhaustively determine if an edge can be identified to connect two subgraphs within a graph using the entire symmetric group of its vertices permutation and via unique permutations we call automorphic subsets, which are a special subset of the symmetric group. We also analyze if edges will be grouped closer to their respective subgraphs in terms of the average algorithmic information contribution. This analysis ascertains if Algorithmic Information Theory can serve as a viable theory for understanding graph substructures and as a foundation for frameworks measuring and analyzing complexity. The study found that the connecting edge was successfully identified as having the highest average information contribution in 29 out of 30 graphs, and in 16 of these, the distance to the next edge was greater than $\log_{2}(2)$. Furthermore, the symmetric group ($S_{G}$) outperformed automorphic subsets in edge grouping.
Author summary
This work reports on the empirical result of the application of the aforementioned complexity values particularly for 2-dimensional binary matrices. By retrieving the complexity of a graph’s matrix using the Block Decomposition Method, we can determine the algorithmic information contribution of an edge before and after it was removed from the graph via a method called edge perturbation. We perform averaging of algorithmic information contribution of edges in graph with respect to different permutations. Since the symmetric group of a graph with $n$ nodes grows by $n!$, runtime also becomes long and thus graphs between 9 and 12 nodes only were analyzed.
Introduction
Complexity has been a topic of interest in science since the 1970s and has sprouted a field called Complex Systems. The field studies the properties and patterns found in a collection of entities interacting with one another within a system [1]. One way of analyzing or understanding complex systems is through networks. Networks are ubiquitous in nature, such as animal social networks, citation networks and social networks [2]. They provide us with a relational view of the structure of interaction of entities in a complex system. Although we use networks to analyze complex systems, the very notion of a complex system is not concrete and no universal definition has been agreed upon in the scientific community.
Finding a clear-cut definition of what constitutes a complex system has been disputed for years now [3]. By nature of complex systems, finding a compressed and simple description of how these systems behave is a difficult task. Since one of the primary objectives of science is to understand nature and reality using unified and simple explanations, defining complexity is embedded in this objective and involves a zoomed out lens view on explaining nature rather than a reductionist approach. By creating a definition of what a complex system is, we can properly analyze these systems that can be found in nature.
In our current scientific climate, to explain phenomenon in nature, we employ the use of computational methods. When we have observational data of a phenomenon, we attempt to find equations / algorithms that explain the phenomenon accurately. We consider our computable functions as explanations of why nature behaves in certain ways. Thus, computable functions– any function that can be executed by a Universal Turing Machine [4], are our compressed descriptions of understanding nature.
One of the candidates of defining complexity while simultaneously aims to provide a computational explanation of observed data which involves the shortest possible description is in the field of Algorithmic Information Theory (AIT). Through the approximations of Kolmogorov-Chaitin complexity, advances in this field have led to the development of the Block Decomposition Method (BDM) [5]. This method computes the complexity of binary strings [6] and networks (via binary adjacency matrices) [7] in a more algorithmic and agnostic perspective rooted from choosing the shortest possible computable function that explains the input data.
Kolmogorov-Chaitin complexity is ultimately uncomputable due to the fact that this function finds the shortest possible computer program that outputs the data and thus can be reduced to the undecidability of the Halting Problem. Consequently, the Block Decomposition Method is an approximation of the algorithmic complexity of binary strings and adjacency matrices. Although they are approximations, it was shown that this method can be used to remove edges from a graph to identify substructures  [8] via an algorithm named causal deconvolution, and it was also shown that this method was used on different networks such as dynamical systems reprogram them move away or towards randomness [9] and analysis of ecological networks [18].
We extend and validate the causal deconvolution algorithm implementation done in [8] by using permutations of vertices to relabel them and then perform edge perturbation. Although the causal deconvolution algorithm managed to identify edges that connect multiple subgraphs such as $K$-ary trees, small-world connected to a complete graphand a graph consisting of a random graph, a complete graph and a star graph, this was only done for one permutation of the graphs’ adjacency matrices which led to the motivation for this work of employing combinatorics to have a more thorough application of BDM to the graphs.
We use the canonical labelling algorithm created by McKay and Piperno [10] to group different permutations from the symmetric group of a graph’s vertices into their respective automorphic subsets. Selected permutations per automorphic subset is used to get the average information contribution of each edge to determine if they will group with other edges that are within the same sub-graph. Comparison is also made when using the entire symmetric group in getting the average information contribution to show which set of permutations is more effective in grouping edges.
Materials and methods
Algorithmic Information Theory
Algorithmic Information Theory uses the program size of Turing machines to determine the randomness of information. It states that, for a given string, if the size of the shortest program that outputs the string is greater than or equal to the string itself, then the string is considered random. This theory formalizes the notion of the complexity of information.
Algorithmic Complexity
The chief equation to this theory is the Kolmogorov-Chaitin $K_{T}(s)$ complexity [11, 12] which is defined as:
$K_{T}(s)=\min\{|p|,T(p)=s\}$ (1)
where it takes a string $s$ and finds the program $p$ that is executable in a (prefix-free) universal Turing machine $T$, and outputs an integer which is the size of the smallest program that outputs $s$. Although this complexity measure has been proven to be uncomputable, approximations have been computed to get the complexity of strings and binary matrices for practical use [6, 7].
Algorithmic Probability
Closely related to algorithmic complexity is the notion of Solomonoff algorithmic probability. This probability determines how likely a random program will output a piece of information [13]. Algorithmic probability is defined as:
$\mathfrak{m}(s)=\sum_{p:T(p)=s}2^{-|p|}$ (2)
where the probability $\mathfrak{m}(s)$ is the sum of probabilities of all (prefix-free) programs $p$ that are executed by a universal Turing machine $T$ and halt. It can be noted that the shorter the length of a program $|p|$, the larger its contribution to the overall summation of program lengths.
Coding Theorem Method
$K_{T}(s)$ can be approximated using $\mathfrak{m}(s)$ by the smallest program in the summation that generates $s$ using the universal Turing machine $T$. The Levin Coding Theorem [14, 15] shows that algorithmic probability and algorithmic complexity are related:
$K(s)\leq\log\frac{1}{\mathfrak{m}(s)}+c$ (3)
where $c$ is a constant that is independent of string $s$. This theorem shows that the complexity of a string is inversely proportional to its algorithmic probability, meaning that the more probable a string is to be generated by a random program, the lower its complexity.
This has theorem been applied to create the Coding Theorem Method (CTM) [6] by enumerating Turing machines from shorter to longer ones based on the Busy Beaver function [16]. Thus, we can approximate $K(s)$ from $\mathfrak{m}(s)$ using the output distribution of halting Turing machines:
$D(t,k)(s)=\frac{|\{T\in(n,m):T(p)=s\}|}{|\{T\in(n,m):T\text{ halts}\}|}$ (4)
We can assign a probability value to a string $s$ by running the Busy Beaver function for $t$ states with $k$ symbols (0,1 for binary sequences). Instead of counting the number of 11s produced by a Turing machine, we count the number of Turing machines that produce $s$ over the number of all Turing machines for $t$ states and $k$ symbols that halt.
With this method, we can use $D(t,k)(s)$ as an approximation of $\mathfrak{m}(s)$ and compute algorithmic complexity:
$CTM(s,t,k)=-\log D(t,k)(s)$ (5)
Block Decomposition Method
Since CTM is based on the Busy Beaver function, and the latter being the fastest growing function compared to all computable functions, CTM is computationally expensive and ultimately non-computable if to be performed. To circumvent this problem, BDM was introduced [5] to extend the capabilities of CTM:
$BDM(s,l,m)=\sum_{i}CTM(s^{i},t,k)+\log(n_{i})$ (6)
We now fix the values of $t$ and $k$ to have a specific distribution table based on $D(t,k)$, which is represented as CTMCTM based in Equation 5. Let $l$ be an integer such that $l\leq s$ and is the length of each substring $s^{i}$ when $s$ is decomposed (with a possible remainder of $y=s\mod l$ characters). After decomposition, there is a probability that some $s^{i}$ substrings are the same, thus $n_{i}$ is the multiplicity of each $s^{i}$ found. If $p_{i}$ is the smallest program that produces substring $s^{i}$, then $\log(n_{i})$ is the number of bits needed to encode $p_{i}$.
Finally, $m$ is an overlapping parameter to handle possible remainder $y$ when $s$ is not a multiple of $l$. If $m=l$, then the remaining characters in the string are ignored. If $m<l$, then a sliding window of size $l$ is used and moves $m$ characters until all parts of the string are captured by the window.
Block Decomposition Method for 2D matrices
CTM has also been used in two-dimensional matrices via two-dimensional Turing machines [7]. Instead of a one-dimensional tape, the Turing machine’s memory is on a two-dimensional grid where up and down is also a possible movement of the Turing machine. As a consequence, BDM can also be extended for graphs using their adjacency matrices and is defined as:
$BDM_{2D}(X,l)=\sum_{(x_{i},n_{i})\in X_{(l\times l)}}CTM_{2D}(x_{i})+\log(n_{i})$ (7)
where $X$ is an $m\times m$ matrix of a graph with $m$ nodes. The sub-matrix size $l$, where $l\leq m$, is used to partition $X$ by $l\times l$ sub-matrices distinctly named $x_{i}$ per sub-matrix being encountered $n_{i}$ times. Currently, 2D CTM values that were computed were up to $4\times 4$, which will be used in the algorithm. If $l$ is not a multiple of $m$, there will be $r$ rows and $c$ columns that are not captured by the partitions, handling of these rows and columns will be also discussed in the algorithm implementation.
Graph Theory
A graph $G$ consists of $(V,E)$ where $V_{G}$ is a set of $n$ vertices labeled $\{1,2,\ldots,n\}$ and $E_{G}$ is the set of edges $\{(u,v)\mid u,v\in V_{G}\}$ that connect vertices. A graph $G$ with $n$ vertices has a symmetric group $S_{G}$ which is the set of all permutations of $V$. Each permutation $\sigma\in S_{G}$, maps a vertex label to another vertex label, $\sigma:V\rightarrow V$, where the edges $E^{\sigma}=\{(\sigma(u),\sigma(v))\mid u,v\in V\}$. We can represent the relabeled graph as $G^{\sigma}$. A sub-graph $G_{1}$ of $G$ is a graph which is defined as $V_{G_{1}}\subseteq V_{G}$ and $E_{G_{1}}\subseteq E_{G}$.
An arbitrary graph $H$ is considered isomorphic to $G$ if there exists a function $\pi:E_{G}\rightarrow E_{H}$ such that $(u,v)\in E_{G}$ and $(\pi(u),\pi(v))\in E_{H}$ for all $u,v\in E_{G}$, where $E_{G}$ and $E_{H}$ are edge sets of each graph respectively. We denote $G\simeq H$ to say that both graphs have an isomorphism. Each permutation $\sigma\in S_{G}$ applied to $G$ preserves edge adjacency, thus $G\simeq G^{\sigma}$. Although vertices are relabeled, the structure of the graph is preserved.
An automorphism $\gamma\in S_{G}$ is also an isomorphism of $G$ (with itself) but with a much stricter requirement of $(\gamma(u),\gamma(v))\in E\iff(u,v)\in E$, meaning an automorphism not only preserves graph structure but also edge adjacency per label of a vertex. Given the original labelling of vertices $\{1,2,\ldots,n\}$, there exists a set of permutations $\{\gamma_{1},\gamma_{2},\ldots\}\subseteq S_{G}$ that are automorphisms of $G$ which is called the automorphism group $\mathcal{A}(G)$. We use $G\overset{\mathcal{A}}{\simeq}H$ to say that two graphs $G,H$ are automorphic.
Canonical Labelling
Canonical labelling was introduced in [10] and summarized in [17], it uses vertex degree and automorphisms to find a definitive relabelling of a graph to test isomorphism with another canonically labelled graph. The canonical labelling function $C$ relabels the vertices of a graph $G$ such that if $G$ and $H$ are isomorphic graphs, then $C(G)=C(H)$.
Let $c_{G}$ be the permutation of the vertices of the graph $G$ when it is passed to $C$. Using $G$ as the original graph, we can apply $\gamma,\sigma\in S_{G}$, to produce two permuted graphs namely $G^{\gamma}$ and $G^{\sigma}$. Upon applying $C$ to both permuted graphs, if $c_{G^{\gamma}}=c_{G^{\sigma}}$, it means that $G^{\gamma}\overset{\mathcal{A}}{\simeq}G^{\sigma}$. Since $G^{\gamma},G^{\sigma}$ are just permuted graphs from $G$, having the same $c$ means that although both graph’s nodes are relabeled differently, when $c$ is applied to them both, their vertex labels and edges are the same, making them automorphic. We use $c$ to group different permutations of $G$ so that each $\sigma\in S_{G}$ is included in what we call an automorphic subset.
Graph Automorphic Subsets
We can now define an automorphic subset of $G$ as:
$\lambda=\{\gamma,\sigma,\phi\ldots\ \mid\gamma,\sigma,\phi,\ldots\in S_{G}\land c_{G^{\gamma}}=c_{G^{\sigma}}=c_{G^{\phi}}=\ldots\}$ (8)
Each element in $\lambda$ is a permutation of $V(G)$ such that when $V(G)$ is permuted by any element in this set, it is automorphic to any other permutation in the same set. This ultimately groups each permutation $\gamma\in S_{G}$ to a specific automorphic subset. Since a graph can have a number of automorphic subsets, we define the complete automorphic set of a graph as:
$\Lambda(G)=\{\lambda_{1},\lambda_{2},\lambda_{3},\ldots\}$ (9)
Each member of $\Lambda(G)$ is a set and also a distinct subset of $S_{G}$. Note that combining all sub-members of all the members of $\Lambda(G)$ gives us $S_{G}$. The number of automorphic subsets within a symmetric group can be determined by the following scenarios:
  • 1, if the graph is completely symmetric (a complete graph), $\{\mathcal{A}(G)\}=\Lambda(G)$.
  • $1<x<|S_{G}|$, where $x$ is the number of automorphic subsets.
  • $|S_{G}|$, if the graph is completely asymmetric (a fully random graph)
Algorithms
If two permutations are automorphic when $G$ is permuted using these permutations, it means that their canonical label permutation $c$ is the same. Thus, it can be used to pick an arbitrary permutation $\gamma$ in $\lambda$ where $\lambda\in\Lambda(G)$. Pynauty [19] - a Python implementation of the canonical labelling package Nauty [10] was used in Algorithm 1 which retrieves unique permutations in $S_{G}$ so that $G$ can be structurally represented distinctly per automorphic subset. We only choose one member from each subset $\lambda\in\Lambda$ since each member of $\lambda$ is already automorphic with every other member of the subset thus, one member can represent the entire subset. It is to note that
$\Cup(G)=\bigcup_{i=1}^{|\Lambda(G)|}\lambda_{i}[x]$ (10)
where $i$ is the $i$th member of $\Lambda(G)$ and $[x]$ retrieves an arbitrary member of $\lambda_{i}$.
Algorithm 1 Automorphic Subset Sampling
[…]
  $\lambda\leftarrow\{\}$}
  $L\leftarrow\{\}$}
  for each $\gamma\in S_{G}$ do
   if $c_{G^{\gamma}}\not\in L$ then
     $\lambda\leftarrow\lambda\cup\{\gamma\}$}
     $L\leftarrow L\cup\{c(G^{\gamma})\}$}
   end if
  end for
  $\textbf{return }U$}
end procedure
Algorithm 2 Average Information Loss via Automorphic Subsets
procedure $\mathfrak{C}$($G$, $\Gamma$)
  $I\leftarrow\{\}$}
  $l\leftarrow 3$}
  $\triangleright$ Initialize a set of tuples where the first member of a tuple is an edge of $G$
  for each $(u,v)\in E_{G}$ do
   $I[(u,v)]\leftarrow 0$}
  end for
  for each $\gamma\in\Gamma$ do
   $X\leftarrow Adj(E_{G^{\gamma}})$}
   $\text{bdm}_{X}\leftarrow BDM_{2D}(X,l)$}
   $\triangleright$ This loop is where edge perturbation happens
   for each $(w,x)\in E_{G^{\gamma}}$ do
     $E_{r}\leftarrow E_{G^{\gamma}}\;\backslash\;(w,x)$}
     $Y\leftarrow Adj(E_{r})$}
     $\text{bdm}_{Y}\leftarrow BDM_{2D}(Y,l)$}
     $\text{bdm}_{(w,x)}\leftarrow\text{bdm}_{X}-\text{bdm}_{Y}$}
     $(u,v)\leftarrow(\gamma^{-}(w),\gamma^{-}(x))$}
     $I[(u,v)]=I[(u,v)]+\text{bdm}_{(w,x)}$}
   end for
  end for
  $\triangleright$ Average all the info loss values using the size of the set of permutations
  for each $(u,v)\in I$ do
   $I[(u,v)]\leftarrow I[(u,v)]\;/\;|\Gamma|$}
  end for
  $\triangleright$ Sort the edges from highest to lowest average info loss
  I = sort(I)
  $\textbf{return }I$}
end procedure
Edge perturbation [9] is the process of temporarily removing an edge from a graph to determine how much algorithmic information is loss or gained upon removal. To know the information contribution of each edge, we get the difference of $BDM_{2D}$ values before and after an edge is removed from a graph. Losing information implies that the graph is moving towards simplicity while gaining information implies it’s moving towards randomness. Since an increase in information implies that a longer Turing machine is needed to output the graph without the edge compared to when it had it. With this method, we can try and group edges using their information contribution values on the basis that $BDM_{2D}$ measures the complexity of a graph. We used PyBDM [20], the Python package that implements $BDM_{2D}$.
$\mathfrak{C}$ (Algorithm 2) uses a set of permutations $\Gamma$ to permute the vertices of a graph $G$ and perform edge perturbation. Initially, a set of tuples II is created where the average algorithmic information contribution of each edge $(u,v)\in E_{G}$ will be housed and can be retrieved by using $(u,v)$ as a key in II which is denoted by $I[(u,v)]$. Using a permutation $\gamma\in\Cup(G)$, we permute $G$ and retrieve its adjacency matrix $X$ that is denoted by $Adj(E_{G^{\gamma}})$. The row and column of the matrix are arranged according to the ordered labelling using $V_{G^{\gamma}}$, where each $v\in V_{G^{\gamma}}$ is the $v$th row/column in the matrix $X$.
For each permuted edge $(w,x)=(\gamma(u),\gamma(v))$ where $u,v\in V_{G}$, we compute its information contribution value $\text{bdm}_{(w,x)}$ by removing $(w,x)$ from $G^{\gamma}$ which is denoted by $E_{G^{\gamma}}\;\backslash\;(w,x)$. After removal, the difference in $BDM_{2D}$ between the original matrix $X$ and the matrix $Y$ (where the edge is removed) is added to the existing value stored in II. To properly add the information contribution for the edges when permuted by $\gamma$, we revert $(w,x)$ to its original label in $E_{G}$ by the inverse function $\gamma^{-}$ since calculating the information contribution of each edge is done in each permutation $\gamma\in\Cup(G)$. Finally, each information contribution value of an edge in II is averaged by the number of $\gamma\in\Cup(G)$ since $|\Cup(G)|=|\Lambda(G)|$.
Recalling the definition of $BDM_{2D}$(7), $l$ is the size of the 2D $l\times l$ sub-matrices produced when $X$ is partitioned. Based on initial testing, it was found that $l=3$ proved to be more effective in grouping edges compared to $l=4$. We conjecture that since the graphs (9 - 12 nodes) analyzed are small, having smaller partitioned sub-matrices is more favorable. We chose the periodic partitioning when converting a matrix $X$ into blocks of $l\times l$ sub-matrices. This partitioning uses the first $|X|\mod l$ rows/columns as padding when the adjacency matrix is not divisible by $l=3$.
Overall, averaging the information contribution of each edge was done to probabilistically account for every possible unique vertex relabelling. This allows for a more holistic accounting of the algorithmic information contribution for each edge to the structure of a graph.
Time Complexity
The algorithm $\mathfrak{C}$ has a runtime of $\mathcal{O}(n!)$ regardless if $\Gamma=\Cup(G)$ or $\Gamma=S_{G}$ since $\Cup(G)$ iterates over $S_{G}$ when sampling a permutation for each automorphic subset $\lambda\in\Lambda(G)$ for a graph $G$. Inside $\mathfrak{C}$, the edge perturbation algorithm is executed and has a linear running time of $\mathcal{O}(m)$, where $m$ is the number of edges in $G$.
Results
Edge perturbation was performed on 30 synthetic graphs where each contains two subgraphs connected by a single edge. We shall label the edge with the highest average information contribution $\max_{\text{info}}$} for any of the 30 graphs. For probabilistically generated graphs like an Erdős–Rényi graph, a Barabási–Albert graph or a Watts–Strogatz graph, we used the same graphs for different vertex counts (for each type of graph respectively) to have a more controlled and rigid experiment of testing edge perturbation. The mentioned graphs have the following properties set:
  1. 1.
    Erdős–Rényi - $p=0.5$, where $p$ is the probability of an edge forming between two vertices.
  2. 2.
    Barabási–Albert - $m=2$, where $m$ is the number of sampled nodes where a newly added node will connect to.
  3. 3.
    Watts–Strogatz - $p=0.5$ & $k=4$, where $p$ is the probability that an edge will rewire and $k$ is the degree of each vertex.
We introduce some more notation on the algorithmic complexity of graphs with subgraphs before the discussion of results. Let $G$ be a graph that has two subgraphs, $G_{1}$ and $G_{2}$ that are connected by a single edge. Then the algorithmic complexity of $G$ is
$K(G)\leq K(G_{1})+K(G_{2})+O(1)$ (11)
where $O(1)$ accounts for the constant program size to connect two vertices from each sub-graph by an edge whose program size is independent of $G$. Let $G^{*}$ be the shortest program that produces $G$ such that $K(G)=|G^{*}|$. If $G$ has two subgraphs, then $|G^{*}|\leq|G^{*}_{1}|+|G^{*}_{2}|+O(1)$. When computing for $K(G)$, if edges $e_{i},e_{j}\in E_{G_{1}}$, then the bits needed to encode $e_{i}$ and $e_{j}$ into $G$ are included in $G_{1}^{*}$. There are three main modes of inquiry that can be asked to determine the effectivity of edge perturbation using BDMBDM:
  • Does the connecting edge for each graph have a high information contribution value when compared to other edges in the same graph?
  • Do edges that belong to the same sub-graph have similar information contribution values?
  • Does the connecting edge have a far enough difference from the other edges that it can be identified as being produced by a different sub-program in $G$?
The following three sections will further expound on these questions respectively. We compared the average information values when using an element in each subset of $\Lambda(G)$ against the symmetric group $S_{G}$ of a graph $G$. In practice, computing for the automorphic subsets $\Lambda(G)$ is usually faster (especially for regular graphs) than using $S_{G}$ since $\Cup(G)$ reduces the number of permutations that need to be edge perturbed unlike for $S_{G}$ that considers all permutations of a $G$ which grows $n!$ for $n$ vertices.
Average Information Contribution of Edges
We checked if the connecting edge’s information contribution is aligned with theoretical expectations in terms of having a high positive information contribution to a graph. If the information contribution of an edge is positive, it implies that information will be lost when the edge is removed. Conversely, if an edge’s information contribution is negative, it means that the algorithmic complexity of the graph increased when the edge was removed. The connecting edge links the two subgraphs together and thus should contribute greatly to the algorithmic complexity of the graph. Viewing the connecting edge as a causal entity, meaning that this edge can only be created once both nodes of which it is connected to are created, then it should be that the connecting edge should not only have a positive information contribution but should also contribute highly to the algorithmic complexity of the graph. Table 1 shows the effectiveness of edge perturbation via algorithm $\mathfrak{C}$. For both $\Gamma=\Cup(G)$ and $\Gamma=S_{G}$ in each graph $G$, $\mathfrak{C}(G,\Gamma)$ was mostly able to identify the connecting edge as the edge with the highest average positive information contribution.
When using $S_{G}$ as the parameter for $\mathfrak{C}$, the algorithm was able to identify that the connecting edge had the highest average information contribution value for all but one graph compared to $\Cup(G)$ (rows colored red in Table 1). $\Cup(G)$ is already effective in identifying if the connecting edge is $\max_{\text{info}}$} but using $S_{G}$ was better because it identified the connecting edge as $\max_{\text{info}}$} for more graphs compared to $\Cup(G)$. The single graph where the algorithm was not successful is the graph where the subgraphs were two Erdős–Rényi graphs with 5 vertices each. This should be correct because for most cases of random graphs, if two substructures are algorithmically random, then connecting them by a single edge is not distinguishable with writing a program that lists all the edges where the connecting edge is also included.
Table 1: A list of synthetic graphs where each graph has two subgraphs connected by a single edge.
Connected Graphs ($G$) $|V_{G}|$ $|E_{G}|$ $|\Lambda(G)|$ $|\Lambda(G)|$ / $|S_{G}|$ Using $\Lambda(G)$ Using $S_{G}$
Complete4 - Cycle5 9 12 30,240 9%
Complete5 - Cycle4 9 15 7,560 3%
Complete4 - Random5 9 13 30,240 9%
Complete5 - Cycle5 10 16 75,600 3%
Complete5 - Star5 10 15 25,200 1%
Complete5 - Complete5 10 21 3,150 1%
Star5 - Random5 10 11 604,800 17%
Random5 - Random5 10 13 1,814,400 50%
Cycle5 - Star5 10 10 302,400 9%
Cycle4 - Star6 10 10 75,600 3%
Cycle5 - Ladder6 11 13 19,958,400 50%
Cycle5 - Random6 11 15 19,958,400 50%
Watts-Strogatz6 - Cycle5 11 18 3,326,400 9%
Complete5 - Random6 11 20 1,663,200 5%
Cycle5 - Star6 11 11 831,600 3%
Watts-Strogatz6 - Complete5 11 23 277,200 1%
Watts-Strogatz6 - Star5 11 17 1,108,800 3%
Barabási–Albert7 - Complete4 11 17 3,326,400 9%
Barabási–Albert7 - Cycle4 11 15 9,979,200 25%
Barabási–Albert6 - Random5 11 15 19,958,400 50%
Random6 - Random5 11 16 39,916,800 100%
Complete6 - Complete5 11 26 13,860 1%
Ladder6 - Random5 11 14 19,958,400 50%
Ladder6 - Complete5 11 18 831,600 3%
Ladder6 - Star5 11 12 3,326,400 9%
Watts-Strogatz6 - Random6 12 22 79,833,600 17%
Watts-Strogatz7 - Complete5 12 25 19,958,400 5%
Watts-Strogatz7 - Star5 12 19 79,833,600 17%
Barabási–Albert7 - Star5 12 15 39,916,800 9%
Complete6 - Complete6 12 31 16,632 1%
The first column shows the names of the two subgraphs that were connected and how many vertices they have, e.g. Cycle5 - Star6 means a cycle graph with 5 vertices and a star graph with 6 vertices. The second and third column shows the number of vertices $V_{G}$, number of edges $E_{G}$ of each graph respectively. The forth column displays how many automorphic subsets $|\Lambda(G)|$ each graph has, and the fifth column is the portion of the amount of automorphic subsets within its symmetric group. The higher the percentage, the more random the graph is, since having many automorphic subsets means the graph is not compressible to a limited number of permutations. The last two columns shows a check mark if $\mathfrak{C}$ was able to determine if the connecting edge has the highest average information contribution in graph $G$ using: (a) one permutation for each automorphic subset, $\mathfrak{C}(G,\Cup(G))$ (sixth column) and (b) all permutations from the entire symmetric group $\mathfrak{C}(G,S_{G})$ (last column). The red rows show the graphs whose edge with the highest average information contribution is not the connecting edge when using one permutation from each automorphic subset (marked with an x mark in the sixth column).
Edge Grouping
To further determine the effectiveness of edge perturbation, we also analyzed how similar the information contribution of an edge to other edges. Since we have computed the average information contribution for every edge of a graph $G$ using $\mathfrak{C}$, we can then compare if edges per sub-graph ($G_{1}$ & $G_{2}$) have grouped correctly using the sorted set of tuples II (recall that II is arranged from highest to lowest in average information contribution value). When $\mathfrak{C}$ outputs II, it shows the likelihood of each edge to be grouped with other edges that were produced by the same underlying sub-program. We categorize each graph into three different grouping schemes with respect to the linear arrangement of the average information contribution values of $E_{G}$:
  1. 1.
    Complete - all the edges in II have grouped $G_{1}$ and $G_{2}$’s edges respectively.
  2. 2.
    Partial - 60% or more of either $G_{1}$ or $G_{2}$’s edges grouped together.
  3. 3.
    Scattered - no edges of either $G_{1}$ or $G_{2}$ were grouped together with at least 60% from the same sub-graph.
Examples of these grouping schemes can be seen in Figure S1 Fig..
Averaging of information contribution values has been done because different vertex permutations result in different $K(G)$ values. Since $\Lambda(G)$ groups isomorphisms of different vertex permutations that are automorphic per grouping, $\Cup(G)$ is the compressed version of $\Lambda(G)$ and ultimately of $S_{G}$. Pairing this approach of getting unique permutations and averaging of each edge’s information contribution in a graph allows us to have a fair assessment of BDMBDM by probabilistic and combinatorial means. We computed the average information contribution values with respect to the parameter $\Gamma$ in $\mathfrak{C}(G,\Gamma)$ to determine if $\Cup(G)$ is already effective without using $S_{G}$ since $\Cup(G)$ is algorithmically quicker to execute. Tables 2 and 3 show a comparison of $\Cup(G)$ and $S_{G}$. We can see that the symmetric group is more effective compared to using automorphic subset sampling. Although using $\Cup(G)$ has partially grouped edges of some graphs, $S_{G}$ correctly grouped edges for more graphs.
Table 2: Edge grouping schemes using an element in each automorphic subset of each graph.
$\Gamma=\Cup(G)$
Complete Partial Scattered
Complete5 - Cycle4 Barabási–Albert7 - Cycle4 Barabási–Albert6 - Random5
Watts-Strogatz7 - Complete5 Barabási–Albert7 - Star5 Barabási–Albert7 - Complete4
Complete4 - Cycle5 Complete5 - Complete5
Complete4 - Random5 Complete6 - Complete5
Complete5 - Cycle5 Complete6 - Complete6
Complete5 - Random6 Cycle5 - Ladder6
Complete5 - Star5 Cycle5 - Star5
Cycle4 - Star6 Ladder6 - Random5
Cycle5 - Random6 Ladder6 - Star5
Cycle5 - Star6 Random5 - Random5
Ladder6 - Complete5 Star5 - Random5
Random6 - Random5 Watts-Strogatz6 - Complete5
Watts-Strogatz6 - Cycle5
Watts-Strogatz6 - Random6
Watts-Strogatz6 - Star5
Watts-Strogatz7 - Star5
Table 3: Edge grouping schemes using the symmetric group of each graph.
$\Gamma=S_{G}$
Complete Partial Scattered
Barabási–Albert7 - Complete4 Barabási–Albert7 - Star5 Barabási–Albert6 - Random5
Barabási–Albert7 - Cycle4 Complete4 - Cycle5 Complete4 - Random5
Complete5 - Cycle4 Complete5 - Complete5 Cycle5 - Ladder6
Complete5 - Cycle5 Complete6 - Complete5 Random5 - Random5
Complete5 - Random6 Complete6- Complete6 Random6 - Random5
Complete5 - Star5 Cycle4 - Star6
Cycle5 - Star6 Cycle5 - Random6
Ladder6 - Complete5 Cycle5 - Star5
Ladder6 - Star5 Ladder6 - Random5
Watts-Strogatz7 - Complete5 Star5 - Random5
Watts-Strogatz6 - Complete5
Watts-Strogatz6 - Cycle5
Watts-Strogatz6 - Random6
Watts-Strogatz6 - Star5
Watts-Strogatz7 - Star5
Partial grouping has been achieved, but with a more refined grouping where at least two edges of a sub-graph cluster together to form a group.
A side-by-side comparison in Figures 1 to 5 highlights further the effectiveness of using $S_{G}$. To highlight, we can also observe that some partially grouped edges in some graphs have shown improvements in terms of their linear arrangement when $\Gamma=S_{G}$ over $\Gamma=\Cup(G)$. These graphs are marked with in Table 3 and their improved information contribution grouping can be seen in Figures 4 and  5.
Fig 1: Graphs whose initial edges grouping was scattered when using $\Gamma=\Cup(G)$(left) but was classified as complete when $\Gamma=S_{G}$(right).
Fig 2: Graphs whose initial edges grouping was partial when using $\Gamma=\Cup(G)$(left) but was classified as complete when $\Gamma=S_{G}$(right).
Fig 3: Graphs whose initial edges grouping was partial when using $\Gamma=\Cup(G)$(left) but was classified as complete when $\Gamma=S_{G}$(right).
Fig 4: Graphs that are categorized as scattered or partial edge grouping when $\Gamma=\Cup(G)$(left) but was refined to have a better grouping when $\Gamma=S_{G}$(right) was used.
Fig 5: Graphs that are categorized as scattered or partial edge grouping when $\Gamma=\Cup(G)$(left) but was refined to have a better grouping when $\Gamma=S_{G}$(right) was used.
Information Contribution Distance
Finally, edge perturbation not only tries to determine how much information does an edge contribute to the overall complexity of a graph but also shows the causal relation of each edge to the entire graph. Since the connecting edge is causally dependent on the existence of $G_{1}$ and $G_{2}$ of a graph $G$, it is theoretically expected to have a distance greater than $\log_{2}(2)$ from the other edges for most graphs (assuming it is not a totally random graph). In Equation 11, the connecting edge $e^{\prime}$ is included in $O(1)$. If $v_{G_{1}}$ and $v_{G_{2}}$ are the vertices (each belong to each sub-graph) that are connected by $e^{\prime}$, then $G_{1}^{*}$ and $G_{2}^{*}$ should be executed / instantiated first so that $e^{\prime}$ can connect the two subgraphs $G_{1}$ and $G_{2}$. The presence of $e^{\prime}$ increases the algorithmic complexity of the graph in terms of not only program size, but also causal relations from previously executed sub-programs when instantiating the entirety of $G$.
Using a combination of the information contribution of an edge and the growth of program lengths, we can determine which edges are instantiated by which sub-programs in $G^{*}$ (the shortest program that produces $G$). In general, the growth of program lengths is $\log_{2}(2)$ [8], this implies that when performing edge perturbation on edge $e_{i}$ and $e_{j}$ (for example), where both edges are included in the same sub-graph, their information contribution values should have a distance of not more than $\log_{2}(2)$ from one another. This is because that there are exponentially shorter programs to describe a program than there are to describe longer ones, thus it is exponentially unlikely that two edges with a difference of $\log_{2}(2)$ will be generated by the same program.
Table 4: The information contribution distance of the connecting edge for each of the 30 graphs tested.
Connected Graphs ($G$) $e_{n}-e_{n-1}$
$\Gamma=\Cup(G)$ $\Gamma=S_{G}$
Barabasi-Albert6 - Random5 0.13378793460520733 0.13378793486272755
Barabasi-Albert7 - Complete4 1.3376391875417948 1.502790069936819
Barabasi-Albert7 - Cycle4 0.5571016664781094 1.0879645196663352
Barabasi-Albert7 - Star5 0.5638933790503007 0.5638933794518453
Complete4 - Cycle5 2.8631256102071503 3.172163330027856
Complete4 - Random5 1.3977474661420275 1.3977474661462583
Complete5 - Complete5 21.6912793586853 23.317882074217472
Complete5 - Cycle4 5.317601412837774 6.002081335753412
Complete5 - Cycle5 2.595326258259141 4.183456312163415
Complete5 - Random6 3.857727731304253 3.857727730900918
Complete5 - Star5 2.792869861977472 4.512949749130205
Complete6 - Complete5 15.485835510254873 17.817978351324065
Complete6 - Complete6 23.27621702970346 24.880830306329372
Cycle4 - Star6 0.4198132053221766 3.6647074566079105
Cycle5 - Ladder6 0.06675395441072052 0.19422595474726378
Cycle5 - Random6 0.04686230316845297 0.04686230326476615
Cycle5 - Star5 0.25489223766848035 1.4481659979161563
Cycle5 - Star6 0.06969126690328409 1.2786870792130305
Ladder6 - Complete5 3.0520261513245543 3.099060700177053
Ladder6 - Random5 0.39179993216664855 0.4665223320742333
Ladder6 - Star5 0.34964872859838714 0.8657191703469174
Random5 - Random5 0.027091365649044796 0.14635893537649114
Random6 - Random5 0.22022158873730113 0.22022158867519082
Star5 - Random5 0.5541784484738965 0.5541784484676509
Watts-Strogatz6 - Complete5 7.936125911627891 7.936125911614662
Watts-Strogatz6 - Cycle5 1.5279233787798026 1.6305962768312012
Watts-Strogatz6 - Random6 2.6902761749130186 2.690276175861212
Watts-Strogatz6 - Star5 1.4985545797543267 1.5750285631331034
Watts-Strogatz7 - Complete5 4.317408673610543 4.317408676679262
Watts-Strogatz7 - Star5 0.9944753991092607 0.9944753983040977
The connecting edge does not have $\max_{\text{info}}$ using this permutation set. Let $e_{n}$ be the connecting edge which is placed in the $n$th location in $I_{k}=\{e_{n+1},e_{n},e_{n-1},\ldots,e_{1}\}$, where $I_{k}$ is the set of keys for the set of tuples II and $I[e_{n}]$ is the average information contribution value of the connecting edge. If the connecting edge is $\max_{\text{info}}$, then $n=|I|$ (the first value in the linear arrangement).
As seen in Table 4, not only did $\mathfrak{C}$ identify that the connecting edge was $\max_{\text{info}}$} for most graphs using both $\Gamma=\Cup(G)$ and $\Gamma=S_{G}$, but it also managed to identify the distance of the connecting edge to the next edge in the linear arrangement of average information contribution was more than the cutoff for half of the graphs. Out of the 30 graphs, 16 of them satisfied $e_{n}-e_{n-1}>\log_{2}(2)$. Interestingly, it shows that performing averaging of information contribution with either $\Cup(G)$ or $S_{G}$ was effective when graphs contain a sub-graph that has a high edge count. Among these 16 graphs, either a Watts-Strogatz or a Complete graph was present as a sub-graph for each graph. This indicates that BDMBDM is effective in detecting tightly connected substructures within a graph. Further proof is that when a graph only contains these two subgraphs, the connecting edge’s average information contribution is high relative to the other graphs examined.
Conversely, BDMBDM did not satisfy $e_{n}-e_{n-1}>\log_{2}(2)$ for 14 out of the 30 graphs (colored red in Table 4) when $\Gamma=\Cup(G)$. 6 of the 9 graphs have a random sub-graph that is included per graph. When both subgraphs are random graphs, then adding a connecting edge will not be distinguishable from the other edges in the graph because the entire graph would be considered one combined random graph. There was not much change in $e_{n}-e_{n-1}$ when using $S_{G}$ over $\Cup(G)$ except for graphs that had a cycle sub-graph. This indicates that averaging of information contribution using $S_{G}$ provides a more fine-grained inspection of substructures. Appending to the findings in the Edge Grouping section of the results, most of the graphs that were classified as complete or partial edge grouping consist of mostly graphs with regular and small-world subgraphs, namely Cycle, Complete and Watts-Strogatz graphs.
Discussion
There are 2 limitations that prohibit us from further experimentation of larger graphs. First is that $\mathfrak{C}$ grows by $n!$, where $n$ is the number of nodes of the graph. To fully test BDMBDM on the basis of combinatorial and probabilistic means, much bigger graphs should be tested using $\mathfrak{C}$. Although $\mathfrak{C}$ shows promise of identifying substructures in a graph and was able to:
  • Determine that the connecting edge was $\max_{\text{info}}$} except for 1 graph using $S_{G}$
  • Completely (10 graphs) or partially (15 graphs) group edges based on their average information contribution in a linear arrangement using $S_{G}$
  • Differentiate the connecting edge by having a distance of more than $\log_{2}(2)$ from the other edges in 20 graphs when using $S_{G}$
it must be tested in real world networks whose number of nodes is greater than 12 to have a more accurate testing of BDMBDM. The second limitation is in the sub-matrix size that CTMCTM (within BDMBDM) can slice an adjacency matrix of a graph. Currently, BDMBDM can handle $3\times 3$ or $4\times 4$ sub matrices and uses the periodic partitioning when $n$ is not divisible by $l$. Ideally, since BDMBDM is an approximated upper-bound of algorithmic complexity, we would like for CTMCTM to handle $l>4$ for larger graphs since there is an error overhead whenever $n$ is not divisible by $l$ and thus the remaining rows and columns in the adjacency matrix are appended with the previous rows and columns using a partitioning technique. But this is a monumental task since computing for larger matrices in CTMCTM is equivalent to running the Busy Beaver function.
On a tangential topic of recursively applying edge perturbation. We can determine if an edge can be cut from a graph (to reveal substructures) when an edge fits the two criteria that we have discussed in the results, which are:
  • the edge needs to have a high information contribution compared to the other edges
  • the edge should have a distance of at least $\log_{2}(2)$
If these criteria have been satisfied by an edge, then we can remove it from the graph. The graphs that have satisfied these criteria are the 16 graphs in Table 4 (rows not colored red). By removing edges in the graph we can recursively apply edge perturbation to the new graph since the adjacency matrix of the new graph has been altered and thus different information contribution values will be assigned to the remaining edges. Using the new information contribution values, we can determine which edges can be removed along with edge grouping (using the linear arrangement of these values) to identify substructures. The more connecting edges are removed, the more edge grouping becomes accurate. For example, if a graph $G$ has three subgraphs namely, $G_{1},G_{2},G_{3}$. If there are edges that connect the subgraphs with one another, then they are removed using edge perturbation recursively until they have all been removed, the algorithmic complexity of $K(G)\leq|G_{1}^{*}|+|G_{2}^{*}|+|G_{3}^{*}|$. This allows $\mathfrak{C}$ to approximate the edges that belong to the subgraphs (which are sub-programs of $G$) only and no added sub-programs are needed since there are no remaining connecting edges. If $m$ is the number of recursions we apply edge perturbation and $n$ is the number of nodes in the graph, then the runtime of $\mathfrak{C}$ with respect to $m$ is $mn!$. Although this is still $\mathcal{O}(n!)$, computing $\mathfrak{C}$ becomes more expensive when recursion is applied.
Conclusion
A combination of permuting a graph then averaging the information contribution per edge perturbed was done to 30 synthetic graphs to determine the efficacy of BDMBDM as a method in revealing substructures within complex networks. The sets of permutations that the algorithm $\mathfrak{C}$ (which uses BDMBDM) used was the automorphic subsets $\Cup(G)$ of a graph and its symmetric group $S_{G}$. We used these two permutation sets to compare a compressed representation of a graph’s structure (automorphic subsets) from its entire permutation set (symmetric group). Out of the 30 graphs, 29 of them have been successfully identified as having the connecting edge as the edge with the highest average information contribution. Within these 29 graphs, 16 were also identified of having the connecting edge as the edge with a distance of more than $\log_{2}(2)$ from other edges within their respective graphs. To highlight, edge perturbation was proven to be effective for graphs that have tightly connected subgraphs, namely Watts-Strogatz graphs and Complete graphs. The connecting edge was clearly differentiated in terms of its average information contribution from the edges of the subgraphs. Averaging of information contribution using $\Cup(G)$ and $S_{G}$ was also effective in completely or partially grouping edges together. Comparing the two permutation sets, $S_{G}$ performed better since more graphs’ edge grouping has classified as a complete edge grouping.
There are several improvements / suggestions we offer to extend experimentation with this methodology. One avenue is to add more edges when connecting one or more subgraphs. We have only connected the two subgraphs by one edge to have a simple testing scenario. By adding more edges or increasing the number of subgraphs, we can have varying results since the graphs are more diverse and may also mimic real-life networks. A second avenue is instead of using automorphic subsets or the symmetric group of the graph, a random sample of permutations with a fixed size of $10!$ can be used so that regardless of the size of the graph, we can have constant time in testing a high enough number of graphs whose number of nodes is greater than 12. And a last avenue is how we assign information contribution values to an edge. In this paper, we have used averaging as a means of assignment, other possible assignment options can be the standard deviation of the information contribution of a perturbed edge from the average algorithmic complexity of $S_{G}$ or a smaller permutation set. Another option is getting the minimum algorithmic complexity in $S_{G}$ then relating that to a perturbed edge’s information contribution when its graph $G$ is permuted by $\sigma\in S_{G}$.
Supporting information
S1 Fig.
Group scheming for graphs (using $\Gamma=\Cup(G)$). First: Complete and Cycle graph (Complete), Second: Watt-Strogatz and Cycle (Partial), Last: Barabási–Albert and Erdős–Rényi (Scattered).
Acknowledgments
I would like to acknowledge my thesis adviser Briane Samson for guiding and motivating me when writing this paper. I would also like to thank Hector Zenil for answering my inquiries regarding their initial implementation of causal deconvolution along with Allan Zea, who implemented the R package of causal deconvolution. Lastly, I would like to show my gratitude to the maintainer of the Python package that implements BDMBDM, Szymon Talaga, for answering questions regarding implementing edge perturbation to the package.